OpenStudy (helpmeplease14):

A company manufactured 1,000 televisions. Testing showed that 20 of the televisions were defective. PART A: What is the experimental probability that the next television willl be defective? PART B: Based on the probability in PART A, how many of the next 5,000 televisions manufactured should the company expect to be defective?

OpenStudy (helpmeplease14):

don't know how to start it off

Miracrown (miracrown):

The prob of a defective TV is the number of defective divided by the number of TV's tested

Miracrown (miracrown):

\[a) \space Prob \space (defective) \space = \frac{ 20 }{ 1000 } \space = \frac{ 2 }{ 100 } \space = \frac{ 1 }{ 50 } \space = .02 \]

Miracrown (miracrown):

\[b) \space Expected \space number \space defective \space = .02(5000) \space = 100\]

OpenStudy (helpmeplease14):

Miracrown

Miracrown (miracrown):

Yes?

Miracrown (miracrown):

The expected number of defectives was 100 - Does it make sense? :)

OpenStudy (helpmeplease14):

yeah thx!!!!!!!! :)

Miracrown (miracrown):

Yw :)